A constant force is applied to a 10 kg box, giving it an acceleration of 25 m/s2.

  1. What is the magnitude of the applied force?
  2. How much work has been done when the box moves 100 m?
Answer #1

Criven: m = 10 kg a= 25 m/s2 a) Fa ma = (10kg ) *(250/52) = 250N b) W = F. d = 250N) * (100m) = 25000 J